OpenStudy (sarahc):

Write the sum using summation notation, assuming the suggested pattern continues. 729 + 1000 + 1331 + 1728 + ... + n3

OpenStudy (anonymous):

let's see! what number is we cubed it (or multiplied by itself 3 times) we get 729

OpenStudy (anonymous):

Can you think of any number?

OpenStudy (anonymous):

that is the first step! we need to find where the summation starts from

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

or let's do something else look at the next term which is 1000 what number ( )^3=1000

OpenStudy (anonymous):

what number give us 1000 if we cubed it?

OpenStudy (anonymous):

this one should be easy

OpenStudy (anonymous):

also 1331 kinda easy! no

OpenStudy (anonymous):

Hey! engage in discussion please! otherwise i get bored talking to myslef

OpenStudy (anonymous):

here some thing useful to do when you don't know factor of a number so 729=3^6=9^3

OpenStudy (anonymous):

Prime factorization is always helpful to factor big composite numbers

OpenStudy (anonymous):

now then we have 9^3+10^3+11^3+12^3+.....+n^3 because 1000=10^3 and 1331=11^3 and you can make sure that 12^3=1728 so that's our summation we need to go further

OpenStudy (aum):

\[\large 729 + 1000 + 1331 + 1728 + ... + n^3 = \\ \large 9^3 + 10^3 + 11^3 + 12^3 +....+n^3 = \\ \large \sum_{k = 9}^{n}k^3 \]

OpenStudy (anonymous):

yes like @aum did our series start from k=9 and continues to n

OpenStudy (anonymous):

I wanted you to work this out by yourself though lol